OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -5 f(1) = 20 f(n) = -4*f(n -1) – 3*f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200, …

OpenStudy (perl):

$$ f(0)=-5 \\ f(1)=20 \\ f(2)=-4f(1) -3*f(0) $$

OpenStudy (perl):

$$ f(0)=-5 \\ f(1)=20 \\ f(2)=-4f(1) -3*f(0)= -4(20) - 3*({-5}) $$

OpenStudy (anonymous):

?

OpenStudy (anonymous):

crap those question marks are there

OpenStudy (perl):

does that make sense

OpenStudy (anonymous):

Sorry, the question marks are not intentional, just a problem with format

OpenStudy (perl):

ok :)

OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -5 f(1) = 20 f(n) = -4*f(n -1) - 3*f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -5, -20, -65, -200,ect -5, 20, -92, 372,ect -5, -24, -92, -372, ect -5, 20, -65, 200, ect